OpenStudy (anonymous):

the distance around an indoor running track is 1/6 mi aruna jogged around the track 16 times how far did she jog

OpenStudy (anonymous):

16 x (1/6) = 16/6 =8/3 =2 and 2/3 miles

OpenStudy (anonymous):